OpenStudy (anonymous):

Write the equation in standard form. State whether the graph of the equation is a parabola, circle, ellipse, or hyperbola . x2 + y2 -8x -6y +5 = 0 a. Ellipse b. Circle c. Hyperbola d. Parabola

OpenStudy (anonymous):

9x^2-36x+36=4y^2+24y+72 Complete square on the x-terms and on the y-terms separately: 9(x^2-4x+4) = 4(y^2+6y+9) + 36 Divide thru by 36 to get: [(x-2)^2/4] - [(y-3)^2/9] = 1